OpenStudy (anonymous):

How do i find all of the zeros for the equation: 3x^3-x^2-15x+5=0 by dividing?!

myininaya (myininaya):

i think would be best to factor by grouping x^2(3x-1)-5(3x-1)=0 (x^2-5)(3x-1)=0

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

So, the x's would be plus of minus 5, amd 1/3?